The recursive definition of a function X is given as:
f(0)=5 and f(n)=f(n-2)+5
Now, find out the value of f(14) using the above function.
1
Expert's answer
2020-10-13T17:42:16-0400

"f(0)=5"

"f(2)=f(0)+5=5+5=10"

"f(4)=15"

If n is even then "f(n)=5*(n\/2 + 1)"

"f(14)=5*(14\/2 + 1)=5 * 8=40"